OpenStudy (anonymous):

50 3 over 4 divided by 6 1 over 2??

OpenStudy (anonymous):

can you change mixed fractions to improper fractions??

OpenStudy (anonymous):

yes i got 200 over 3 times 13 over 2

OpenStudy (anonymous):

are you sure that is what you got??

OpenStudy (anonymous):

you are not supposed to get 200/3.......try it again

OpenStudy (anonymous):

203 over 4

OpenStudy (anonymous):

ok...what you need to know is that \[\frac{ \frac{ a }{ b } }{ \frac{ c }{ d } }=\frac{ a }{ b } \times \frac{ d }{ c }\]

OpenStudy (anonymous):

can you carry on from there??

OpenStudy (anonymous):

Not really

OpenStudy (anonymous):

using what i gave above..you now have \[\frac{ 203 }{ 4 } \times \frac{ 2 }{ 13 }=??\]

OpenStudy (anonymous):

do you understand it now or a little??